Zeronknight (zeronknight):

Find the unit rate for the ratio of 1/2 to 1/4 * 4:1 2:1 1/2 : 1 1/4 : 1

Zeronknight (zeronknight):

@Mehek14

Mehek (mehek14):

\(\bf\dfrac{1}{2} \div \dfrac{1}{4}\) \(\bf\dfrac{1}{2}x\dfrac{4}{1}\)

Zeronknight (zeronknight):

1 x 4 = 4 2 x 1 = 2 4/2. 2/4?

Mehek (mehek14):

4/2 = ?

Zeronknight (zeronknight):

4/2 = I don't even know?

Mehek (mehek14):

\(\bf\dfrac{1}{2}x\dfrac{4}{1}=\dfrac{4}{2}=2\)

Zeronknight (zeronknight):

OH

Mehek (mehek14):

so 2:1

Zeronknight (zeronknight):

2:1?

Zeronknight (zeronknight):

xD

Mehek (mehek14):

yea